OpenStudy (anonymous):

How do I express my question's answer, ab^2/1/a^2b^3 using only positive exponents?

OpenStudy (anonymous):



OpenStudy (anonymous):

ab^2/(1/a^2b^3) =(ab^2)(a^2b^3) =a^3 b^5 but if it is just what you say that ab^2/(a^1/2 b^1/3) =a^(1- 1/2) b^(2- 1/3) =a^1/2 b^5/3